# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v1

@@ Abstract
whose mixed tensor rank equals one. Using a general geometric bound for connections with skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
@@ 1 Introduction
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the compact product manifold.
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
@@ 3.3 Vanishing of the obstruction kernel $\mathcal{K}$
Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus
@@ 4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
The physical metric (2.1) is not a Riemannian product, whereas Theorem 5.2 of PT [12] is stated for genuine product metrics. The following three results show that the theorem nevertheless applies to our Kaluza–Klein geometry, because the relevant topological and parallel-form data coincide with those of an auxiliary product metric.
@@ 4.1 Parallel forms in Kaluza–Klein metrics
Lemma 4.1 (Parallel forms in Kaluza–Klein metrics with flat fibres).
@@
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$ (4.1)
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$ (possibly with $dA^{a}\neq 0$).
Then the restriction map
$\Phi_{1}:\mathcal{P}_{k}(M_{1},g)\longrightarrow\mathcal{P}_{k}(M_{1},g_{1}),\qquad\eta\mapsto\eta|_{D_{1}},$
is a well-defined linear isomorphism.
Analogously, define $\mathcal{P}_{k}(M_{2},g)$ as the space of forms $\xi\in\Omega^{k}(M)$ that are invariant along $M_{1}$ (their coefficients depend only on the $M_{2}$-coordinates) and satisfy $\nabla^{LC}\xi=0$. Then
$\Phi_{2}:\mathcal{P}_{k}(M_{2},g)\longrightarrow\mathcal{P}_{k}(M_{2},h),\qquad\xi\mapsto\xi|_{D_{2}},$
is an isomorphism.
@@
Choose local coordinates $(x^{i},y^{\alpha})$ adapted to the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the product structure. In these coordinates the metric reads
$g=g_{ij}(x)\,dx^{i}dx^{j}+h_{\alpha\beta}(y)\bigl(dy^{\alpha}+A^{\alpha}_{i}(x)dx^{i}\bigr)\bigl(dy^{\beta}+A^{\beta}_{j}(x)dx^{j}\bigr).$ (4.2)
Let $\eta\in\mathcal{P}_{k}(M_{1},g)$. Because $\eta$ is parallel and tangent to $D_{1}$, in the coordinates above it can be written as
$\eta=\frac{1}{k!}\,\eta_{i_{1}\dots i_{k}}(x)\,dx^{i_{1}}\wedge\dots\wedge dx^{i_{k}}.$
We compute the Christoffel symbols of $g$. From (4.2) one obtains
$\Gamma^{i}_{jk}$ $=\Gamma^{(g_{1})i}_{jk},$
$\Gamma^{i}_{j\beta}$ $=-\frac{1}{2}\,g^{i\ell}\bigl(\partial_{j}g_{\ell\beta}+\partial_{\beta}g_{j\ell}-\partial_{\ell}g_{j\beta}\bigr),$
$\Gamma^{\alpha}_{jk}$ $=\frac{1}{2}\,h^{\alpha\gamma}\bigl(\partial_{j}g_{k\gamma}+\partial_{k}g_{j\gamma}-\partial_{\gamma}g_{jk}\bigr),$
$\Gamma^{\alpha}_{\beta\gamma}$ $=\Gamma^{(h)\alpha}_{\beta\gamma},$
where $\Gamma^{(g_{1})}$ and $\Gamma^{(h)}$ denote the Christoffel symbols of $g_{1}$ and $h$ respectively.
For any vector field $X=X^{i}\partial_{i}+X^{\alpha}\partial_{\alpha}$,
$(\nabla^{LC}_{X}\eta)_{i_{1}\dots i_{k}}=X^{j}\partial_{j}\eta_{i_{1}\dots i_{k}}-\sum_{r=1}^{k}\Gamma^{\ell}_{ji_{r}}X^{j}\eta_{i_{1}\dots i_{r-1}\ell i_{r+1}\dots i_{k}}-\Gamma^{\alpha}_{j\alpha}X^{j}\eta_{i_{1}\dots i_{k}}.$
The condition $\nabla^{LC}\eta=0$ therefore splits into two independent sets of equations. For $X=X^{i}\partial_{i}$ (tangent to $M_{1}$) we obtain
$\nabla^{(g_{1})}_{X}\eta=0,$
where $\nabla^{(g_{1})}$ is the Levi‑Civita connection of $g_{1}$. Hence $\eta|_{D_{1}}$ is $g_{1}$-parallel.
For $X=X^{\alpha}\partial_{\alpha}$ (tangent to $M_{2}$) we get
$\Gamma^{\alpha}_{j\alpha}(x)X^{j}\eta_{i_{1}\dots i_{k}}=0\qquad\text{for all }X^{j}.$
Why $\Gamma^{\alpha}_{j\alpha}=0$. Because the fibres are flat, the fibre metric $h_{\alpha\beta}$ is constant in the chosen coordinates; therefore $\partial_{j}h_{\alpha\beta}=0$. Moreover, from (4.2) we have $g_{j\alpha}=h_{\alpha\beta}A^{\beta}_{j}(x)$, which is independent of $y$. Consequently
$\partial_{\gamma}g_{j\alpha}=0,\qquad\partial_{\alpha}g_{j\gamma}=0.$
Substituting these facts into the expression for $\Gamma^{\alpha}_{j\alpha}$ gives
$\Gamma^{\alpha}_{j\alpha}=\frac{1}{2}h^{\alpha\gamma}\bigl(\partial_{j}h_{\alpha\gamma}+\partial_{\alpha}g_{j\gamma}-\partial_{\gamma}g_{j\alpha}\bigr)=0.$
Thus the condition coming from $X^{\alpha}\partial_{\alpha}$ is automatically satisfied.
Hence $\eta\in\mathcal{P}_{k}(M_{1},g)$ if and only if $\eta|_{D_{1}}\in\mathcal{P}_{k}(M_{1},g_{1})$. Consequently $\Phi_{1}$ is injective. Surjectivity follows because any $g_{1}$-parallel form $\widetilde{\eta}$ on $M_{1}$ can be lifted to a $g$-parallel form on $M$ by declaring it independent of the fibre coordinates; the computation above shows that the lifted form is indeed parallel with respect to $g$.
Therefore $\Phi_{1}$ is an isomorphism. The proof for $\Phi_{2}$ proceeds identically after exchanging the roles of base and fibre. ∎
Lemma 4.2 (Harmonic Künneth decomposition for flat‑fibre KK metrics).
Let $g$ be a Kaluza–Klein metric as in (4.1) with flat fibres. Then the natural wedge‑product map
$\Phi:\mathcal{H}^{p}(M_{1},g_{1})\otimes\mathcal{H}^{q}(M_{2},h)\longrightarrow\mathcal{H}^{p+q}(M,g)$
is an isomorphism onto the subspace of harmonic forms of pure bidegree $(p,q)$. In particular, every $g$-harmonic form $\eta\in\mathcal{H}^{p+q}(M,g)$ that belongs to the $(p,q)$-Künneth component admits a unique decomposition
$\eta=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i},\qquad\alpha_{i}\in\mathcal{H}^{p}(M_{1},g_{1}),\;\beta_{i}\in\mathcal{H}^{q}(M_{2},h),$
with $r=\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}})$.
@@
Because the fibres are flat, the Hodge Laplacian of $g$ decouples as $\Delta_{g}=\Delta_{g_{1}}\otimes\mathrm{Id}+\mathrm{Id}\otimes\Delta_{h}$ on forms of pure bidegree. Hence a form of bidegree $(p,q)$ is $g$-harmonic iff its components are harmonic with respect to $g_{1}$ and $h$ respectively. The statement then follows from the classical Künneth theorem for de Rham cohomology combined with Hodge theory. ∎
@@ 4.2 Persistence of non‑parallelism under metric change
For the application of the PT theorem to an auxiliary product metric, we need to ensure that the non‑parallelism condition $\nabla^{LC}T\neq 0$ is preserved when we replace the original Kaluza–Klein metric by a nearby product metric.
@@
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly (in fact, rationally) on the metric $g$ and its first derivatives. Consequently, the tensor field $\nabla^{LC}(g)T$ depends continuously on $g$ in the $C^{1}$-topology. The condition $\nabla^{LC}(g)T\neq 0$ means that at least one component of this tensor is non‑zero at some point; by continuity, the same component remains non‑zero for all metrics sufficiently close to $g$ in the $C^{1}$-sense. ∎
@@
In our setting, the original metric $g$ is the Kaluza–Klein metric (4.1) and the auxiliary product metric $g_{\mathrm{prod}}=g_{1}\oplus h$ can be chosen arbitrarily close to $g$ in the $C^{\infty}$ (hence also $C^{1}$) topology by taking the gauge potentials $A^{a}$ small, or by a direct approximation argument. Lemma 4.3 guarantees that if $\nabla^{LC}(g)T\neq 0$, then also $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ for a generic choice of $g_{\mathrm{prod}}$ near $g$.

### v2

@@ Abstract
whose mixed tensor rank equals one. Using a general geometric bound for metric connections with totally skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
@@ 1 Introduction
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the manifold
$M=T_{BZ}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
We emphasize that treating $M$ as a global product manifold requires the underlying principal bundle to be trivial, which holds precisely in the physical regime of vanishing total Chern number ($c_{\text{tot}}=0$) that is central to our main result. In configurations with non-zero net topological charge, this product structure should be understood as a local description in a gauge patch, sufficient for deriving local curvature bounds.
@@ 3.3 Vanishing of the obstruction kernel $\mathcal{K}$
By Lemma 4.5 (with $A^{(\pm)}$ non‑flat), the space of parallel 1‑forms on the phase circles vanishes: $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus
@@ 4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
4.1 Deformation to a product metric and stability
The physical Kaluza–Klein metric (2.1) is a warped structure that does not admit a global Riemannian product decomposition, whereas Theorem 5.2 in [12] is formally stated for genuine product metrics. To bridge this gap, we introduce a smooth deformation to a product metric and prove the stability of the holonomy bound.
Definition 4.1 (Deformation family).
Let $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ be the smooth one-parameter family of metrics on $M$ defined by:
$g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})^{2}+(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})^{2}.$
For $\varepsilon=0$ we obtain the genuine Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, while for $\varepsilon=1$ we recover the physical Kaluza–Klein metric $g_{M}=g_{1}$.
Lemma 4.2 (Stability of the obstruction data).
For the family $g_{\varepsilon}$ in Definition 4.1, the following hold:
  1. 1.
    The spaces of parallel forms satisfy:
    $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\varepsilon})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{0})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}\quad\text{for all }\varepsilon\in[0,1].$
  2. 2.
    For $\varepsilon>0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{\varepsilon})=\{0\}$. For $\varepsilon=0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{0})=\mathbb{R}d\phi_{\pm}$.
  3. 3.
    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
  4. 4.
    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon\in[0,1]$.
Proof.
(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$. (2) For $\varepsilon>0$, Lemma 2.4 shows $d\phi_{\pm}$ are not parallel; the proof of Lemma 4.5 will show no other parallel 1-forms exist. For $\varepsilon=0$, $d\phi_{\pm}$ are clearly parallel. (3) is Lemma 4.6 (metric invariance).
(4) For $\varepsilon=0$, the harmonic representative of $\omega]$ is
$\omega_{h}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}.$
The subspace $\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$ consists of those harmonic 3‑forms that can be written as
$\alpha\wedge\beta,$
where $\alpha$ is a parallel 2‑form on $T^{2}$ (hence a constant multiple of $\text{vol}_{\text{BZ}}$) and $\beta$ is a parallel 1‑form on the fibres (hence a constant linear combination of $d\phi_{+}$ and $d\phi_{-}$).
If $\omega_{h}$ belonged to this subspace, we would have
$F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\alpha\wedge\beta=(\lambda\text{vol}_{\text{BZ}})\wedge(c_{+}d\phi_{+}+c_{-}d\phi_{-}),$
which would imply $F^{(+)}=\lambda c_{+}\text{vol}_{\text{BZ}}$ and $F^{(-)}=\lambda c_{-}\text{vol}_{\text{BZ}}$ as differential forms, not just in cohomology. In a generic SOC texture, $F^{(+)}$ and $F^{(-)}$ are not parallel (they are not constant multiples of $\text{vol}_{\text{BZ}}$ pointwise), hence $\omega_{h}\notin\mathcal{P}_{2}\otimes\mathcal{P}_{1}$. Therefore $\mathcal{K}_{0}=0$. ∎
Lemma 4.3 (Lower semi-continuity of off-diagonal holonomy dimension).
The function $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$, where $\nabla^{C_{\varepsilon}}$ is the metric connection with fixed torsion $T$ constructed from $g_{\varepsilon}$.
Proof.
The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$. At any point $p\in M$, consider the evaluation map
$\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\times T_{p}M\to\mathfrak{so}(T_{p}M),\quad\Phi_{\varepsilon}(X,Y,Z)=R^{C_{\varepsilon}}(X,Y)Z.$
The off-diagonal holonomy algebra at $p$ is generated by the image of $\Phi_{\varepsilon}$ restricted to arguments that yield off-diagonal operators. The dimension of a finitely generated subspace is a lower semi-continuous function of the generating set (it can only jump upward at discontinuity points). Since $\Phi_{\varepsilon}$ varies continuously in $\varepsilon$, $F(\varepsilon)$ is lower semi-continuous. ∎
Proposition 4.4 (Persistence of the PT bound under deformation).
If the Pigazzini–Toda theorem gives $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}$ for the product metric $g_{0}$, then the same inequality holds for all $\varepsilon\in[0,1]$, and in particular for the physical metric $g_{M}=g_{1}$.
Proof.
By Lemma 4.2, the topological data $r^{\sharp}=r-\dim\mathcal{K}_{\varepsilon}$ is constant (=r=r) for all $\varepsilon$. The inequality holds at $\varepsilon=0$ by assumption. By Lemma 4.3, the set $\{\varepsilon:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}\}$ is closed (by lower semi-continuity). Since it contains $\varepsilon=0$ and $[0,1]$ is connected, it must be all of $[0,1]$. ∎
@@ 4.2 Parallel forms in Kaluza–Klein metrics
Lemma 4.5 (Parallel forms in Kaluza–Klein metrics).
@@
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
  1. 1.
    If all connection forms $A^{a}$ are closed (i.e., $dA^{a}=0$), then every $h$-parallel form on the fibre is also $g$-parallel.
  2. 2.
    If at least one curvature $F^{a}:=dA^{a}$ is not identically zero, then the space of $g$-parallel 1-forms tangent to the fibres is trivial:
    $\mathcal{P}_{1}(M_{2},g)=\{0\}.$
@@
Write the metric locally as $g=g_{ij}(x)dx^{i}dx^{j}+\sum_{a=1}^{m}(\Theta^{a})^{2}$, where $\Theta^{a}=\theta^{a}+A^{a}_{i}(x)dx^{i}$ and $\{\theta^{a}\}$ is the orthonormal coframe on $M_{2}$ induced by the flat metric $h$. In our SOC BEC case, $m=2$, $\theta^{1}=d\phi_{+}$, $\theta^{2}=d\phi_{-}$.
Let $\xi=\sum_{a=1}^{m}f_{a}(x,y)\Theta^{a}$ be a 1‑form tangent to the fibres. The Levi‑Civita connection of $g$ satisfies the following relation for a horizontal vector field $X=X^{i}\partial_{i}$ and the vertical coordinate field $V_{a}$ dual to $\theta^{a}$:
$\nabla^{LC}_{X}V_{a}=\tfrac{1}{2}F^{a}(X)^{\#},$
where $F^{a}(X)^{\#}$ is the vertical vector field obtained by raising the index of the 1‑form $F^{a}(X,\cdot)$ with the fibre metric $h$. Formula (i) follows from the Koszul formula and uses only that the fibres are flat.
Now suppose $\xi$ is parallel, $\nabla^{LC}\xi=0$. Evaluating on the pair $(X,V_{a})$ gives
$0=(\nabla^{LC}_{X}\xi)(V_{a})=X\bigl(\xi(V_{a})\bigr)-\xi(\nabla^{LC}_{X}V_{a})=X(f_{a})-\tfrac{1}{2}\sum_{b=1}^{m}f_{b}\,h\!\bigl(F^{a}(X)^{\#},V_{b}\bigr).$
Because the functions $f_{a}$ are constant along the fibres (they depend only on the base coordinates $x$ due to the left‑invariance of the metric $h$), equation (ii) reduces to
$\tfrac{1}{2}\sum_{b=1}^{m}f_{b}(x)\,F^{b}(X,V_{a})=0\qquad\text{for every horizontal }X\text{ and every }a.$
If all $F^{b}=0$, condition (iii) is vacuous and any choice of constant $f_{b}$ yields a parallel form, proving statement (i). If, however, some $F^{b_{0}}$ is not identically zero, we can choose a horizontal vector field $X$ and an index $a$ such that $F^{b_{0}}(X,V_{a})\neq 0$ at some point; then (iii) forces $f_{b_{0}}=0$. Repeating the argument for every independent curvature forces all coefficients $f_{b}$ to vanish, hence $\xi=0$. This proves statement (ii) for 1‑forms.
The argument for higher‑degree forms is analogous, using that a parallel $k$-form on the fibre must have constant coefficients with respect to the left‑invariant coframe, and the same curvature coupling appears in the covariant derivative. ∎
Lemma 4.6 (Harmonic representative and Künneth component).
Let $g$ be a Kaluza–Klein metric as in Definition 4.1 on $M=M_{1}\times M_{2}$. Let $\omega]\in H^{k}(M;\mathbb{R})$ be a cohomology class and $\omega_{h}$ its unique $g$-harmonic representative. Then $\omega_{h}$ respects the Künneth decomposition in the sense that its projection onto the $(p,q)$-bidegree component corresponds to the harmonic representative of the mixed class $\omega]_{p,q}\in H^{p}(M_{1})\otimes H^{q}(M_{2})$. In particular, $\omega_{h}$ admits a unique decomposition
$\omega_{h}=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i}+\gamma,$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$, the forms $\alpha_{i},\beta_{i}$ are the lifted harmonic representatives of the factor classes, and $\gamma$ contains terms of different bidegrees.
@@
By Hodge theory on compact Riemannian manifolds, every de Rham cohomology class $\omega]$ possesses a unique harmonic representative $\omega_{h}$ with respect to the metric $g$. The Künneth theorem provides a canonical isomorphism $H^{k}(M)\cong\bigoplus_{p+q=k}H^{p}(M_{1})\otimes H^{q}(M_{2})$.
While the Hodge Laplacian $\Delta_{g}$ of a Kaluza-Klein metric does not generally commute with the factor-wise Laplacians $\Delta_{g_{1}}$ and $\Delta_{h}$ due to the presence of the connection forms $A^{a}$, the existence of the harmonic representative for each mixed component is guaranteed by the ellipticity of $\Delta_{g}$. The tensor rank $r$ is a purely algebraic property of the class $\omega]$ in the tensor product of the cohomology groups $H^{*}(M_{i})$ and is therefore independent of the specific metric $g$.
The decomposition follows by projecting the global harmonic form $\omega_{h}$ onto the subspace of forms of pure bidegree, which remains well-defined under the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the Kaluza-Klein structure. ∎
@@ 4.3 Persistence of non‑parallelism under metric change
To ensure that the lower bound on the off-diagonal holonomy is a robust feature of the model rather than a consequence of a specific metric choice, we must examine the stability of the non-parallelism condition $\nabla^{LC}T\neq 0$. The following result establishes that this property is open in the space of Riemannian metrics, guaranteeing that the topological obstruction remains effective under generic deformations of the Kaluza–Klein geometry.
@@
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly on the components of the metric $g$ and its first derivatives through the Christoffel symbols. Consequently, the map $g\mapsto\nabla^{LC}(g)T$ is continuous with respect to the $C^{1}$-topology on $\mathcal{M}$. Since the condition $\nabla^{LC}(g)T\neq 0$ defines the complement of the zero set of a continuous map, $\mathcal{M}_{\neq 0}(T)$ is an open subset of $\mathcal{M}$. ∎
@@
In the context of spin-orbit coupled BECs, Lemma 4.7 ensures the structural stability of our topological bound. Since the Kaluza–Klein metric $g$ satisfies $\nabla^{LC}(g)T\neq 0$ due to the non-triviality of the synthetic gauge curvature, any sufficiently small perturbation of the metric—arising, for instance, from experimental fluctuations in the trapping potentials or the SOC strength—will preserve the non-parallelism of the torsion. This confirms that the lower bound on the off-diagonal holonomy is not a fine-tuned property but a robust feature of the SOC BEC geometry.